OpenStudy (anonymous):

Help? A 1150 Kg wrecking ball is released from a height 13.0 meters above its resting point and hits a wall at the lowest point of its swing. If the ball makes a .5 meter indentation in the wall, what is the average force the ball applies to the wall? Assume that the ball uses all of its mechanical energy to do work on the wall.

OpenStudy (anonymous):

146510 N 1150 N 293020 N

OpenStudy (irishboy123):

Kinetic energy (T) of ball just before collision: T = mg∂h = 1150(9.8)(13) K is convert to work (W) done smashing the wall on collision: W = F•∂x = F(0.5), where F = average force Thus F = (2)(1150)(9.8)(13) = ?!?! Do it and see!!